OpenStudy (aye_my_peeps):

WHAT IS THE PERCENT OF INCREASE FROM 60 TO 75? A.80% B.25% C.20% D.15%

OpenStudy (aye_my_peeps):

@Awolflover1

OpenStudy (aye_my_peeps):

@truwhovian11

OpenStudy (aye_my_peeps):

@pooja195

OpenStudy (studygurl14):

Take 75 minus 60, then take that and divide by 60. Convert to % and you're done

OpenStudy (studygurl14):

PS. It's not d

OpenStudy (helasi95):

So, 75 - 60 =15/60 * 100= 25% @StudyGurl14